# Beyond Freezing: Sparse Tuning Enhances Plasticity in Continual Learning with Pre-Trained Models

### v1

@@ Abstract
Continual Learning with Pre-trained Models holds great promise for efficient adaptation across sequential tasks. However, most existing approaches freeze PTMs and rely on auxiliary modules like prompts or adapters, limiting model plasticity and leading to suboptimal generalization when facing significant distribution shifts. While full fine-tuning can improve adaptability, it risks disrupting crucial pre-trained knowledge. In this paper, we propose Mutual Information-guided Sparse Tuning (MIST), a plug-and-play method that selectively updates a small subset of PTM parameters, less than 5%, based on sensitivity to mutual information objectives. MIST enables effective task-specific adaptation while preserving generalization. To further reduce interference, we introduce strong sparsity regularization by randomly dropping gradients during tuning, resulting in fewer than 0.5% of parameters being updated per step. Applied before standard freeze-based methods, MIST consistently boosts performance across diverse continual learning benchmarks. Experiments show that integrating our method into multiple baselines yields significant performance gains. Our code is available at https://github.com/zhwhu/MIST.
1 Introduction
@@
Continual Learning (CL) Chaudhry et al. (2019); Zhou et al. (2024); Lyu et al. (2021, 2023); Liu et al. (2023) is a paradigm in which tasks are learned sequentially, aiming to reduce catastrophic forgetting of previously acquired knowledge while integrating new information. Recently, Pre-Trained Models (PTMs)  Han et al. (2021); Chen et al. (2021) have shown potential to enhance learning efficiency in CL tasks. By fine-tuning, PTMs can be easily adapted to various downstream tasks, enabling continual learners to acquire new task-specific knowledge more effectively and improving resilience to catastrophic forgetting Goodfellow et al. (2013). One important challenge of PTMs in CL lies in how to effectively adapt to incremental tasks without harming the generalization ability of PTMs.
@@
A common practice is to freeze the PTM and introduce additional learnable parameters or modules to adapt the frozen PTM to new tasks. These methods can typically be categorized into two types: prompt-based methods and adapter-based methods. Prompt-based methods, such as L2P Wang et al. (2022b) and DualPrompt Wang et al. (2022a) introduce additional learnable prompt pools, which dynamically guide the frozen pre-trained layers to accommodate incremental tasks. Adapter-based methods, such as APER Zhou et al. (2025) and RanPAC McDonnell et al. (2023), adapt the frozen PTM by introducing additional adapters during the initial incremental stage to bridge the domain gap between pre-trained representations and incremental task distributions. In summary, most PTM-based CL methods typically freeze PTMs during incremental learning, relying heavily on the pure pre-trained knowledge for downstream adaptation. However, new task distributions may deviate from the encoded PTM knowledge, the freezing backbone struggles to generalize effectively across all tasks, that is, poor plasticity Zhang et al. (2023a). Since freezing PTMs can reduce model plasticity, it raises the question of why some methods that fine-tune PTMs still achieve suboptimal performance. A possible explanation lies in the fact that heuristic fine-tuning or fully updating all parameters can lead to the loss of crucial parameters, diminishing the effectiveness of PTMs themselves. To avoid this, effective sparse tuning is needed, which selectively updates only a subset of parameters, thus preserving key knowledge within PTMs. The goal of this paper is to propose a sparse update method that strikes a balance between effective adaptation to new tasks and the preservation of generalization in PTM-based CL methods.
@@
Figure 1: (a) Refer to caption indicates learnable parameters, while Refer to caption denotes frozen parameters. MIST leverages MI for pre-adaptation, enabling it to be plugged into other methods. (b) Comparison of different methods in terms of average accuracy, new task accuracy, and additional parameters. Without requiring additional parameters, MIST achieves superior accuracy through MI-guided sparse tuning.
@@
Therefore, how to selectively identify important parameters in PTM-based CL remains a key challenge. To address this, we investigate the underlying behavior of PTMs through a probabilistic analysis. We theoretically and empirically demonstrate that the parameters sensitive to the MI objective can effectively model task-specific knowledge while minimizing disruption to the original knowledge structure of the PTM. Motivated by this, we introduce a simple yet effective plug-and-play method named Mutual Information-guided Sparse Tuning (MIST). Specifically, before training each incremental task with other freeze-based methods, we first determine the sensitivity of each PTM parameter to the MI objective. We then select the top 5% most sensitive parameters for MI-guided tuning , which enables the model to fully adapt to the new task distribution while maximally preserving the structural knowledge encoded in the PTM. During this process, we apply strong regularization by randomly dropping the gradients of 90% of the selected parameters in each mini-batch, thereby updating only 0.5% of the parameters per batch. After this tuning stage, we freeze the PTM and proceed with the original freeze-based method for continual learning. We insert our approach into five representative freeze-based methods and conduct experiments on serval datasets. Results show consistent performance improvements with MIST, particularly on datasets with large distribution shifts from the pretraining domain. For example, SimpleCIL with MIST achieves 17.9% and 15.7% improvements on Split-ImageNet-R and Split-Cars, respectively. The contributions of this paper are summarized as follows:
  1. (1)
@@
    We study PTM-based CL from a probabilistic and information-theoretic perspective, and theoretically and empirically demonstrate, through MI techniques, PTMs can effectively adapt to new tasks by updating only a small subset of parameters.
  2. (2)
@@
    We introduce a simple yet effective plug-and-play method named Mutual Information-guided Sparse Tuning (MIST), which can be integrated into freeze-based methods to provide significant performance improvements.
  3. (3)
@@
    We incorporate MIST into five representative PTM-based CL methods and evaluate them across five benchmark datasets. All methods achieve consistent performance gains after integrating MIST, highlighting its broad applicability and effectiveness. The empirical results clearly demonstrate the superiority of MIST.
2 Related Work
@@
Continual Learning on a Pre-trained Models. Recently, advancements in PTMs and their exceptional performance in adapting to downstream tasks have inspired researchers to investigate how PTMs can be adapted for continual learning across sequential tasks. Prompt-based methods learn continual prompts to provide fixed PTMs with additional instruction. DualPrompt Wang et al. (2022a) combines task-shared and task-specific prompts to achieve an effective balance between adaptability and mitigating forgetting, while CODA-Prompt Smith et al. (2023) leverages contrastive learning-based prompts to enhance the representation learning of PTMs for improved task adaptation. HiDe-PromptWang et al. (2024) optimizes hierarchical components by combining task-specific prompts and representation statistics, enhanced with a contrastive regularization strategy. Adapter-based methods also freeze the PTM and introduce additional lightweight modules for task-specific adaptation. SLCA++Zhang et al. (2024a) sequentially fine-tunes low-rank LoRA matrices with a small learning rate to avoid disrupting the pre-trained features. RanPACMcDonnell et al. (2023) adapts the PTM during the first task to enhance downstream performance, while APER Zhou et al. (2025) further combines the adapted PTM with the original frozen PTM to jointly extract features, aiming to balance generalization and task-specific learning.
@@
Mutual Information in Machine Learning. With the advancement of deep learning, mutual information (MI) has become an important tool for capturing both linear and nonlinear dependencies between variables, supporting tasks such as feature selection, clustering, and model optimization Zhang et al. (2023b); Vinh et al. (2009); Tishby and Zaslavsky (2015). In particular, InfoNCE Oord et al. (2018) has emerged as a widely used lower-bound estimator of MI in representation learning. Building on this, recent works have applied InfoNCE-based MI objectives to continual learning. For example, Guo et al.Guo et al. (2022) used InfoNCE to measure MI between samples to mitigate catastrophic forgetting, while Li et al.Li et al. (2023) maximized MI between outputs of current and previous models for knowledge distillation. In this work, we construct an MI-guided sparse tuning to identify important parameters during incremental fine-tuning in PTM-based CL, enabling more targeted and generalization-preserving updates.
3 Rethinking the use of PTMs in Continual Learning
3.1 Continual Learning with PTMs and the Impact of Freezing PTMs
@@
Given a sequence of tasks with data $\{\mathcal{D}^{1},\mathcal{D}^{2},\dots,\mathcal{D}^{T}\}$, where $\mathcal{D}^{t}=\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ with $n_{t}$ input pair, sample $x$ and its corresponding label $y$. Different tasks are with disjoint label spaces across tasks: $\mathcal{Y}_{i}\cap\mathcal{Y}_{j}=\emptyset$ for $i\neq j$. At the training stage $t$, only the current task dataset $\mathcal{D}^{t}$ is available. The model is denoted as $f_{\theta}$, where $\theta$ is parameters. In PTM-based CL, $\theta$ is initialized from a PTM trained on a large-scale dataset and is typically frozen during adaptation. The model adapts to new tasks by introducing additional parameters, which can take the form of prompts or adapters, depending on the chosen tuning strategy. Despite their structural differences, both methods freeze the backbone and optimize lightweight parameters for efficient adaptation. Freezing the PTM parameters $\theta$ limits adaptability to new tasks, shifting the burden to auxiliary modules like prompts or adapters.
@@
Freezing PTMs in prompt tuning: In prompt-based tuning (the left subfigure in Fig. 1(a)), learnable prompts $\phi$ are prepended or injected into the input embedding, yielding output $p(y\mid x;\theta,\phi)=f_{\theta}(P_{\phi}(x))$. The gradient of the log-likelihood with respect to $\phi$ follows the chain rule:
$\frac{\partial\log p(y\mid x;\theta,\phi)}{\partial\phi}=\frac{1}{p(y\mid x;\theta,\phi)}\cdot\frac{\partial f_{\theta}(P_{\phi}(x))}{\partial x}\cdot\frac{\partial P_{\phi}(x)}{\partial\phi},$ (1)
@@
where $P_{\phi}(x)$ represents the modified input obtained by injecting the learnable prompt $\phi$ into the feature space of $x$. Since $\theta$ is frozen, the Jacobian term ${\partial f_{\theta}}/{\partial x}$ is fixed and reflects the model’s sensitivity to input perturbations. When this Jacobian is close to zero in directions that encode task-specific features, the gradient signal received by $\phi$ is significantly diminished, regardless of its expressive capacity Qiao et al. (2023); Fu et al. (2024); Gao et al. (2023). This severely restricts the effectiveness of prompt-based tuning, especially under distribution shifts where new tasks require directions outside the pre-trained manifold.
@@
Freezing PTMs in adapter-based tuning: Adapter-based tuning (the center subfigure in Fig. 1(a)) inserts trainable adapters $\psi$ into the intermediate layers, resulting in $p(y\mid x;\theta,\psi)=f_{\theta,\psi}(x)$. The gradient of $\psi$ is:
@@
where $f_{\theta,\psi}$ means the backbone modified by inserting adapter modules. From Eq. (2), adapter’s influence must propagate through the remaining frozen layers to affect the output. If the PTM is not responsive to the features injected by adapters, particularly when such features lie outside the pre-trained distribution, then the resulting gradient with respect to $\psi$ is similarly attenuated Qiao et al. (2024); Son et al. (2024); Nowak et al. (2024).
In summary, despite using different mechanisms, both prompt- and adapter-based methods suffer from gradient suppression due to the fixed representational structure of the frozen PTM. The frozen PTM acts as a bottleneck that limits the flow of gradients to newly introduced parameters. This constraint hampers the model’s ability to adapt to novel tasks.
3.2 Continual Tuning on PTMs
To enhance plasticity in PTM-based CL, some works have explored direct fine-tuning. However, studies Kingma and Ba (2014); Zhang et al. (2024a) show that this often results in significant performance drops, particularly under distribution shifts. To analyze this, we begin by examining the gradient of the log-likelihood:
$\frac{\partial\log p(y\mid x;\theta)}{\partial\theta^{i}}=-\frac{1}{p(x,y;\theta)}\cdot\frac{\partial p(x,y;\theta)}{\partial\theta^{i}}+\frac{1}{p(x;\theta)}\cdot\frac{\partial p(x;\theta)}{\partial\theta^{i}},$ (3)
@@ 3.2 Continual Tuning on PTMs
Figure 2: Zero-shot accuracy.
Existing tuning strategies, including full fine-tuning, naive partial fine-tuning, and Fisher-guided partial fine-tuning, share a common limitation: they inadvertently perturb the pre-trained structure by amplifying the term ${\partial p(x;\theta)}/{\partial\theta^{i}}$, thereby severely reducing the generalization of the PTM (more details in Appendix). As illustrated in Fig. 2, this drawback leads to a continual decline in zero-shot accuracy on the Cars dataset as tasks progress, clearly indicating progressive loss of generalization ability. Existing strategies either overfit to new tasks or disrupt pre-trained generalization due to their inability to disentangle task-relevant gradients from those that compromise structural stability. This motivates the need for a more principled tuning strategy that explicitly controls the influence on each gradient component in Eq. (3). This motivates the need for a more principled tuning strategy that explicitly controls the influence on each gradient component in Eq. (3).
4 Method
4.1 Mutual Information Analysis in PTM-based CL
MI Kraskov et al. (2004); Lei et al. (2023); Zhang et al. (2024b) is a fundamental concept in information theory and has been widely adopted in machine learning. By maximizing the MI $I(X;Y)$ between input $X$ and output $Y$, MI explicitly quantifies the statistical dependency between features and labels Guo et al. (2022). Formally, the MI $I(X;Y)$ is defined as:
@@ 4.1 Mutual Information Analysis in PTM-based CL
where $p(y;\theta)$ denotes the prior probabilities of the target classes. Due to the normalization constraint of probability distributions, we have $\sum_{x}p(x;\theta)=1$ and $\sum_{y}p(y;\theta)=1$. Under this constraint, the gradient of the MI with respect to $\theta^{i}$ can be simplified as (more proofs in Appendix):
@@
In summary, MI provides a more stable optimization objective than CE for CL with PTMs. Unlike CE gradients, which include the marginal term $\partial p(x;\theta)/\partial\theta^{i}$ and may disrupt the pre-trained input distribution, MI gradients inherently avoid this due to normalization constraints, preserving feature integrity. Additionally, MI benefits from diverse batches, which better approximate the underlying data distribution and reduce gradient bias. Together, these properties enable MI to strike a more effective balance between plasticity and stability during adaptation. While MI enables a better plasticity–stability trade-off, directly replacing CE for full fine-tuning may still lead to information loss and high computational cost due to large-scale updates. To address this, we next introduce a lightweight MI-based method that selectively tunes a small parameter subset and can be flexibly integrated as a plugin into existing PTM-based CL methods, including both prompt-tuning and adapter-based approaches.
4.2 Mutual Information-guided Sparse Tuning (MIST)-A plug-and-play solution
@@
To efficiently estimate the sensitivity of each parameter $\theta^{i}\in\theta$ to the MI objective, we adopt the MI-based Fisher Information Matrix Chaudhry et al. (2018) as an importance measure. While computing exact gradients over the entire task is computationally intensive, the sample distribution within a task is typically uniform in CL (more proofs in Appendix), enabling a batch-wise approximation:
@@
where $\mathcal{L}_{\mathrm{MI}}^{\mathcal{D}_{t}}$ denotes the MI loss computed over task $\mathcal{D}_{t}$, and $B_{j}$ represents the $j$-th mini-batch sampled from $\mathcal{D}_{t}$. In practice, we identify the top $k\%$ of parameters with the highest $F^{\prime}_{\mathrm{MI}}$ values as the most MI-sensitive parameters, denoted by $\mathcal{M}$:
=$\mathcal{M}=\left\{\theta^{i}\in\theta\mid\operatorname{rank}(F^{\prime}_{\mathrm{MI}}(\theta^{i}))\leq\lfloor k\%\cdot|\theta|\rfloor\right\}.$ (7)
[…]
@@
6:        Generate dropped parameters group $\mathcal{M}^{\prime}$ by dropping $d\%$ parameters in $\mathcal{M}$
7:        Update the parameters in $\mathcal{M}^{\prime}$
@@
With the MI-sensitive parameter subset $\mathcal{M}$ identified, we proceed to the pre-adaptation stage using an MI-based objective to minimize disruption to the pre-trained feature structure. However, computing the exact MI loss is challenging in practice, as both joint and marginal distributions $p(x,y;\theta)$ and $p(x;\theta)$ are typically intractable. Inspired by OCM Guo et al. (2022), we adopt the supervised InfoNCE loss to construct the MI objective:
@@
$A_{i}=-\sum_{y_{k}=y_{i}}\log\frac{g(x_{i},x_{k})\cdot g(x_{i},x_{k}^{\prime})\cdot g(x_{i}^{\prime},x_{k})}{\left(\sum_{j=1}^{|B|}g(x_{i},x_{j})+g(x_{i},x_{j}^{\prime})+g(x_{i}^{\prime},x_{j})\right)^{3}},$ (9)
where $g(x_{i},x_{j}^{\prime})=e^{\frac{f_{\theta}(x_{i})^{T}f_{\theta}(x_{j}^{\prime})}{\tau}}$ is the similarity of two samples, $\tau$ is temperature, $x_{j}^{\prime}$ is an augmentation view of sample $x_{j}$ (more analysis in Appendix). By optimizing Eq. (8), we effectively maximize the MI I(X;Y), thereby modeling $p(x,y;\theta)$ in a task-discriminative manner.
@@
4.3 Plugging MIST into PTM-based Continual Learning: The Algorithm

### v2

# Sparse Tuning Enhances Plasticity in PTM-based Continual Learning

@@ Abstract
Continual learning with Pre-trained Models (PTMs) holds great promise for efficient adaptation across sequential tasks. However, most existing approaches freeze PTMs and rely on auxiliary modules like prompts or adapters, limiting model plasticity and leading to suboptimal generalization when facing significant distribution shifts. While full fine-tuning can improve adaptability, it risks disrupting crucial pre-trained knowledge. In this paper, we propose Mutual Information-guided Sparse Tuning (MIST), a plug-and-play method that selectively updates a small subset of PTM parameters, less than 5%, based on sensitivity to mutual information objectives. MIST enables effective task-specific adaptation while preserving generalization. To further reduce interference, we introduce strong sparsity regularization by randomly dropping gradients during tuning, resulting in fewer than 0.5% of parameters being updated per step. Applied before standard freeze-based methods, MIST consistently boosts performance across diverse continual learning benchmarks. Experiments show that integrating our method into multiple baselines yields significant performance gains. Our code is available at https://github.com/zhwhu/MIST.
Introduction
@@
Continual Learning (CL) (lyu2021multi) is a paradigm in which tasks are learned sequentially, aiming to reduce forgetting of previously acquired knowledge while integrating new information. Recently, Pre-Trained Models (PTMs)  (han2021pre) have shown potential to enhance learning efficiency in CL tasks. By fine-tuning, PTMs can be easily adapted to various downstream tasks, enabling continual learners to acquire new task-specific knowledge more effectively and improving resilience to catastrophic forgetting (sun2022exploring). One important challenge of PTMs in CL lies in how to effectively adapt to incremental tasks without harming the generalization ability of PTMs.
@@
A common practice is to freeze the PTM and introduce additional learnable parameters to adapt the frozen PTM to new tasks. These methods can typically be categorized into two types: prompt-based methods and adapter-based methods. Prompt-based methods, such as L2P (wang2022l2p) and DualPrompt (wang2022dualprompt) introduce additional learnable prompt pools, which dynamically guide the frozen pre-trained layers to accommodate incremental tasks. Adapter-based methods, such as APER (zhou2025revisiting) and RanPAC (mcdonnell2023ranpac), adapt the frozen PTM by introducing additional adapters during the initial incremental stage to bridge the domain gap between pre-trained representations and incremental task distributions. In summary, most PTM-based CL methods typically freeze PTMs during incremental learning, relying heavily on the pure pre-trained knowledge for downstream adaptation. However, new task distributions may deviate from the encoded PTM knowledge, the freezing backbone struggles to generalize effectively across all tasks, that is, poor plasticity (zhang2023slca). Since freezing PTMs can reduce model plasticity, it raises the question of why some methods that fine-tune PTMs still achieve suboptimal performance. A possible explanation lies in the fact that heuristic fine-tuning or fully updating all parameters can lead to the loss of crucial parameters, diminishing the effectiveness of PTMs themselves. To avoid this, effective sparse tuning is needed, which selectively updates only a subset of parameters, thus preserving key knowledge within PTMs. The goal of this paper is to propose a sparse update method that strikes a balance between effective adaptation to new tasks and the preservation of generalization in PTM-based CL methods.
@@
Figure 1: (a) Refer to caption indicates learnable parameters, while Refer to caption denotes frozen parameters. MIST leverages MI for pre-adaptation, enabling it to be plugged into other methods. (b) Comparison of different methods in terms of average accuracy, new task accuracy, and additional parameters. MIST achieves superior accuracy through MI-guided sparse tuning.
@@
Therefore, how to selectively identify important parameters in PTM-based CL remains a key challenge. To address this, we investigate the underlying behavior of PTMs through a probabilistic analysis. We theoretically and empirically demonstrate that the parameters sensitive to the MI objective can effectively model task-specific knowledge while minimizing disruption to the original knowledge structure of the PTM. Motivated by this, we introduce a simple yet effective plug-and-play method named Mutual Information-guided Sparse Tuning (MIST). Specifically, before training each incremental task with other freeze-based methods, we first determine the sensitivity of each PTM parameter to the MI objective. We then select the top 5% most sensitive parameters for MI-guided tuning , which enables the model to fully adapt to the new task distribution while maximally preserving the structural knowledge encoded in the PTM. During this process, we apply strong regularization by randomly dropping the gradients of 90% of the selected parameters in each mini-batch, thereby updating only 0.5% of the parameters per batch. After this tuning stage, we freeze the PTM and proceed with the original freeze-based method for continual learning. We insert our approach into six representative freeze-based methods and conduct experiments on serval datasets. Results show consistent performance improvements with MIST, particularly on datasets with large distribution shifts from the pretraining domain. For example, SimpleCIL with MIST achieves 17.9% and 15.7% improvements on Split-ImageNet-R and Split-Cars, respectively. The contributions of this paper are summarized as follows:
@@
(1) We study PTM-based CL from a probabilistic and information-theoretic perspective, and theoretically and empirically demonstrate, through MI techniques, PTMs can effectively adapt to new tasks by updating only a small subset of parameters.
@@
(2) We introduce a simple yet effective plug-and-play method named Mutual Information-guided Sparse Tuning (MIST), which can be integrated into freeze-based methods to provide significant performance improvements.
@@
(3) We incorporate MIST into six representative PTM-based CL methods and evaluate them across five benchmark datasets. All methods achieve consistent performance gains after integrating MIST, highlighting its broad applicability and effectiveness. The empirical results clearly demonstrate the superiority of MIST.
Related Work
@@
Continual Learning on Pre-trained Models. Recently, advancements in PTMs and their exceptional performance in adapting to downstream tasks have inspired researchers to investigate how PTMs can be adapted for continual learning across sequential tasks. Prompt-based methods learn continual prompts to provide fixed PTMs with additional instruction. DualPrompt (wang2022dualprompt) combines task-shared and task-specific prompts to achieve an effective balance between adaptability and mitigating forgetting, while CODA-Prompt (smith2023coda-prompt) leverages contrastive learning-based prompts to enhance the representation learning of PTMs for improved task adaptation. HiDe-Prompt(hide-prompt) optimizes hierarchical components by combining task-specific prompts and representation statistics, enhanced with a contrastive regularization strategy. Adapter-based methods also freeze the PTM and introduce additional lightweight modules for task-specific adaptation. SLCA++(zhang2024slca++) sequentially fine-tunes low-rank LoRA matrices with a small learning rate to avoid disrupting the pre-trained features. RanPAC(mcdonnell2023ranpac) adapts the PTM during the first task to enhance downstream performance, while APER (zhou2025revisiting) further combines the adapted PTM with the original frozen PTM to jointly extract features, aiming to balance generalization and task-specific learning.
@@
Mutual Information in Machine Learning. With the advancement of deep learning(lyu2024elastic; yin2025beyond; zhang2024constructing; du2024confidence), mutual information (MI) has become an important tool for capturing both linear and nonlinear dependencies between variables, supporting tasks such as feature selection, clustering, and model optimization (zhang2023mfsjmi; vinh2009information; tishby2015dlib). In particular, InfoNCE (oord2018representation) has emerged as a widely used lower-bound estimator of MI in representation learning. Building on this, recent works have applied InfoNCE-based MI objectives to continual learning. For example, Guo et al.(guo2022ocm) used InfoNCE to measure MI between samples to mitigate catastrophic forgetting, while Li et al.(li2023variational) maximized MI between outputs of current and previous models for knowledge distillation. In this work, we construct an MI-guided sparse tuning to identify important parameters during incremental fine-tuning in PTM-based CL, enabling more targeted and generalization-preserving updates.
Rethinking PTMs in Continual Learning
Continual Learning with PTMs and the Impact of Freezing PTMs
Preliminaries. Given a sequence of tasks with data $\{\mathcal{D}^{1},\mathcal{D}^{2},\dots,\mathcal{D}^{T}\}$, where $\mathcal{D}^{t}=\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ with $n_{t}$ input pair, sample $x$ and its corresponding label $y$. Different tasks are with disjoint label spaces across tasks: $\mathcal{Y}_{i}\cap\mathcal{Y}_{j}=\emptyset$ for $i\neq j$. At the training stage $t$, only the current task dataset $\mathcal{D}^{t}$ is available. The model is denoted as $f_{\theta}$, where $\theta$ is parameter.
@@
In PTM-based CL, $\theta$ is initialized from a PTM trained on a large-scale dataset and is typically frozen during adaptation. The model adapts to new tasks by introducing additional parameters, which can take the form of prompts or adapters, depending on the chosen tuning strategy. Despite their structural differences, both methods freeze the backbone and optimize lightweight parameters for efficient adaptation. Freezing the PTM parameters $\theta$ limits adaptability to new tasks, shifting the burden to auxiliary modules like prompts or adapters.
@@
Freezing PTMs in prompt tuning. In prompt-based tuning (the left subfigure in Fig. 1(a)), learnable prompts $\phi$ are prepended or injected into the input embedding, yielding output $p(y\mid x;\theta,\phi)=f_{\theta}(P_{\phi}(x))$. The gradient of the log-likelihood with respect to $\phi$ follows the chain rule:
$\frac{\partial\log p(y\mid x;\theta,\phi)}{\partial\phi}=\frac{1}{p(y\mid x;\theta,\phi)}\cdot\frac{\partial f_{\theta}(P_{\phi}(x))}{\partial x}\cdot\frac{\partial P_{\phi}(x)}{\partial\phi},$
@@
where $P_{\phi}(x)$ represents the modified input obtained by injecting the learnable prompt $\phi$ into the feature space of $x$. Since $\theta$ is frozen, the Jacobian term ${\partial f_{\theta}}/{\partial x}$ is fixed and reflects the model’s sensitivity to input perturbations. When this Jacobian is close to zero in directions that encode task-specific features, the gradient signal received by $\phi$ is significantly diminished, regardless of its expressive capacity (qiao2023prompt2; fu2024prompt4; gao2023prompt). This severely restricts the effectiveness of prompt-based tuning, especially under distribution shifts where new tasks require directions outside the pre-trained manifold.
@@
Freezing PTMs in adapter-based tuning. Adapter-based tuning (the center subfigure in Fig. 1(a)) inserts trainable adapters $\psi$ into the intermediate layers, resulting in $p(y\mid x;\theta,\psi)=f_{\theta,\psi}(x)$. The gradient of $\psi$ is:
@@
where $f_{\theta,\psi}$ means the backbone modified by inserting adapter modules. From Eq. (1), adapter’s influence must propagate through the remaining frozen layers to affect the output. If the PTM is not responsive to the features injected by adapters, particularly when such features lie outside the pre-trained distribution, then the resulting gradient with respect to $\psi$ is similarly attenuated (qiao2024ada1; son2024ada3; nowak2024ada4).
In summary, despite using different mechanisms, both prompt- and adapter-based methods suffer from gradient suppression due to the fixed representational structure of the frozen PTM. The frozen PTM acts as a bottleneck that limits the flow of gradients to newly introduced parameters. This hampers the model’s ability to adapt to novel tasks.
Continual Tuning on PTMs
To enhance plasticity in PTM-based CL, some works have explored direct fine-tuning. However, studies (adam; zhang2024slca++) show that this often results in significant performance drops, particularly under distribution shifts. To analyze this, we begin by examining the gradient of the log-likelihood:
$\frac{\partial\log p(y\mid x;\theta)}{\partial\theta^{i}}=$ $-\frac{1}{p(x,y;\theta)}\cdot\frac{\partial p(x,y;\theta)}{\partial\theta^{i}}$ (2)
$+\frac{1}{p(x;\theta)}\cdot\frac{\partial p(x;\theta)}{\partial\theta^{i}},$
@@
Figure 2: Zero-shot accuracy using different sparse update strategies. After training on each ImageNet-R task, we freeze the PTM and train only on a new classification head on the full Cars dataset. Higher accuracy reflects better generalization capability of the PTM learned from each task.
Existing tuning strategies, including full fine-tuning, naive partial fine-tuning, and Fisher-guided partial fine-tuning, share a common limitation: they inadvertently perturb the pre-trained structure by amplifying the term ${\partial p(x;\theta)}/{\partial\theta^{i}}$, thereby severely reducing the generalization of the PTM. As illustrated in Fig. 2, this drawback leads to a continual decline in zero-shot accuracy on the Cars dataset as tasks progress, clearly indicating progressive loss of generalization ability. Existing strategies either overfit to new tasks or disrupt pre-trained generalization due to their inability to disentangle task-relevant gradients from those that compromise structural stability. This motivates the need for a more principled tuning strategy that explicitly controls the influence on each gradient component in Eq. (2). This motivates the need for a more principled tuning strategy that explicitly controls the influence on each gradient component.
Method
Mutual Information Analysis in PTM-based CL
MI (EstimatingMI; lei2023mutual) is a fundamental concept in information theory and has been widely adopted in machine learning. By maximizing the MI $I(X;Y)$ between input $X$ and output $Y$, MI explicitly quantifies the statistical dependency between features and labels (guo2022ocm). Formally, the MI $I(X;Y)$ is defined as:
@@
where $p(y;\theta)$ denotes the prior probabilities of the target classes. Due to the normalization constraint of probability distributions, we have $\sum_{x}p(x;\theta)=1$ and $\sum_{y}p(y;\theta)=1$. Under this constraint, the gradient of the MI with respect to $\theta^{i}$ can be simplified as:
@@
In summary, MI provides a more stable optimization objective than CE for CL with PTMs. Unlike CE gradients, which include the marginal term $\partial p(x;\theta)/\partial\theta^{i}$ and may disrupt the pre-trained input distribution, MI gradients inherently avoid this due to normalization constraints, preserving feature integrity. Additionally, MI benefits from diverse batches, which better approximate the underlying data distribution and reduce gradient bias. Together, these properties enable MI to strike a more effective balance between plasticity and stability during adaptation. While MI enables a better plasticity–stability trade-off, directly replacing CE for full fine-tuning may still lead to information loss and high computational cost due to large-scale updates. To address this, we next introduce a lightweight MI-based method that selectively tunes a small parameter subset and can be flexibly integrated as a plugin into existing PTM-based CL methods, including prompt-based and adapter-based approaches.
Mutual Information-guided Sparse Tuning (MIST)-A plug-and-play solution
@@
To efficiently estimate the sensitivity of each parameter $\theta^{i}\in\theta$ to the MI objective, we adopt the MI-based Fisher Information Matrix (ewc_0) as an importance measure. While computing exact gradients over the entire task is computationally intensive, the sample distribution within a task is typically uniform in CL, enabling a batch-wise approximation:
@@
where $\mathcal{L}_{\mathrm{MI}}^{\mathcal{D}_{t}}$ is the MI loss computed over task $\mathcal{D}_{t}$, and $B_{j}$ represents the $j$-th mini-batch sampled from $\mathcal{D}_{t}$. In practice, we identify the top $k\%$ of parameters with the highest $F^{\prime}_{\mathrm{MI}}$ values as the most MI-sensitive parameters, denoted by $\mathcal{M}$:
$\mathcal{M}=\left\{\theta^{i}\in\theta\mid\operatorname{rank}(F^{\prime}_{\mathrm{MI}}(\theta^{i}))\leq\lfloor k\%\cdot|\theta|\rfloor\right\}.$ (6)
where $\operatorname{rank}(\cdot)$ denotes the descending order index, and $\lfloor\cdot\rfloor$ denotes the floor function. That is, we select the top $k\%$ parameters with the highest MI-based importance scores. Given that only a small subset of parameters is updated in each batch and that PTMs are typically initialized near an optimal solution (zhang2023slca; zhou2025revisiting), we compute $F^{\prime}_{\mathrm{MI}}$ once at the beginning of each task and reuse $\mathcal{M}$ throughout the pre-adaptation phase to ensure both efficiency and effectiveness.
@@
6:   Generate dropped parameters group $\mathcal{M’}$ by dropping $d\%$ parameters in $\mathcal{M}$
7:   Update the parameters in $\mathcal{M’}$
@@
With the MI-sensitive parameter subset $\mathcal{M}$ identified, we proceed to the pre-adaptation stage using an MI-based objective to minimize disruption to the pre-trained feature structure. However, computing the exact MI loss is challenging in practice, as both joint and marginal distributions $p(x,y;\theta)$ and $p(x;\theta)$ are typically intractable. Inspired by OCM (guo2022ocm), we adopt the supervised InfoNCE loss to construct the MI objective:
@@
$A_{i}=-\sum_{y_{k}=y_{i}}\log\frac{g(x_{i},x_{k})\cdot g(x_{i},x_{k}^{\prime})\cdot g(x_{i}^{\prime},x_{k})}{\left(\sum_{j=1}^{|B|}g(x_{i},x_{j})+g(x_{i},x_{j}^{\prime})+g(x_{i}^{\prime},x_{j})\right)^{3}},$
where $g(x_{i},x_{j}^{\prime})=e^{{f_{\theta}(x_{i})^{T}f_{\theta}(x_{j}^{\prime})}/{\tau}}$ is the similarity of two samples, $\tau$ is temperature, $x_{j}^{\prime}$ is an augmentation view of sample $x_{j}$. By optimizing Eq. (7), we effectively maximize the MI I(X;Y), thereby modeling $p(x,y;\theta)$ in a task-discriminative manner.
@@
Plugging MIST into PTM-based Continual Learning: The Algorithm